OpenStudy (anonymous):

Part 1: Solve the inequality -5x + 8 <= x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

-5x + 8 = x - 4 Is this how its supposed to look like

OpenStudy (anonymous):

-5x + 8 <= x - 4 srry forgot to fix that

OpenStudy (anonymous):

\[-5x + 8 \le x - 4\]

OpenStudy (anonymous):

−5x + 8 ≤ x − 4 First you subtract x from both sides. -x -x ------------------ -4x + 8 ≤ - 4 Then you subtract 8 from both sides. -8 -8 ---------------- -4x ≤ -12 Then divide by -4 /-4 /-4 --------------- x ≤ 3

OpenStudy (anonymous):

ok thanks for part 1

OpenStudy (anonymous):

Sorry I can't help for the second part I am not really that good with graphing. :(